OpenStudy (anonymous):

it took me forever to solve a problem similar to this, you try it: In a recent year, the total sales at store A were $250,000 more than at store B. Total sales at the two stores were $600,000,000. What were the sales at each store.

OpenStudy (lilg132):

oh my bad

OpenStudy (akshay_budhkar):

haste is waste lol :P

OpenStudy (lilg132):

ts = b + a a = b + 250,000 600,000,000 = b + (b + 250,000) 600,000,000 = 2b + 250,000 600,000,000 - 250,000 = 2b 599,750,000 = 2b store b = 299,875,000 600,000,000 - 299,875,000 = store a store a = 300,125,000

OpenStudy (lilg132):

am i right?

OpenStudy (lilg132):

i guess i am ye took about 1 and a half minutes after i noticed it was 600,000,000 instead of 600,000 :p

OpenStudy (akshay_budhkar):

umm yea i feel u r right o.O

OpenStudy (akshay_budhkar):

lol

OpenStudy (lilg132):

thanks :p

OpenStudy (anonymous):

Excuse me as I was sidetracked but unfortunately, your answer was incorrect. It was, B=299,862,500 and A=300,137,500.

OpenStudy (akshay_budhkar):

OMG

OpenStudy (akshay_budhkar):

let me see it again

OpenStudy (anonymous):

I keep doing different problems to see if I have the concept down, but I don't know where I am going wrong.

OpenStudy (akshay_budhkar):

ummm i dun see anything wrong with lilg has done to be truthful fionaJ.. it seems pretty correct :O

OpenStudy (akshay_budhkar):

are u sure the answer is correct?

OpenStudy (akshay_budhkar):

the numbers r big so maybe an error in addition is possible, but the concept is correct.. do u get the concept?

OpenStudy (anonymous):

not really.. "/

OpenStudy (akshay_budhkar):

um ok.. can u see the solution given by lilg the big one.. in that which step did u not comprehend?

OpenStudy (anonymous):

Pardon me, I do understand the concept. I just dnt understand how the answer is always wrong when I follow every step !

OpenStudy (akshay_budhkar):

ok lets solve this together now..let the sales of store A be a and that of store B be B

OpenStudy (akshay_budhkar):

now we have the sales of A where 250000 more than b so the equation is a=b+250000

OpenStudy (akshay_budhkar):

then it is given the total sales of A and B were 6000000 so we have a+b=600,000,000

OpenStudy (akshay_budhkar):

now lets substitute the first equation in second we have b+250000 + b=600,000,000

OpenStudy (akshay_budhkar):

umm you get it till here?

OpenStudy (akshay_budhkar):

fionaJ?

OpenStudy (lilg132):

fionaj unfortunately YOUR answer is incorrect a way to check your answer is remember store a got $250,000 more sales then store b so to check if your answer is correct you add 250,000 to store b's sales and it should be the same as the answer for store a another way to check is to add both store a and b together you should get 600,000,000 as you can see with your answer B=299,862,500 and A=300,137,500. store b = 299,862,500 299,862,500 + 250,000 = 300,112,500 that is not the same answer as you got for store A hence why your answer is incorrect :) dont forget to click good answer

OpenStudy (anonymous):

- excuse me as I had an emergency to tend to, but I have put this dwn into my notes, thank you all so much. I understand it now.

OpenStudy (anonymous):

And I answered this on one of my reviews and your answer, which was MY answer as well, is incorrect.

OpenStudy (lilg132):

I think your slightly confused fionaj your answer that YOU wrote was B=299,862,500 and A=300,137,500. My answer that I wrote was store a = 300,125,000 store b = 299,875,000 what was the correct answer then if you say my answer was incorrect?

OpenStudy (anonymous):

My answer was simply what you had but the computer generated review was B=299,862,500 and A=300,137,500.

OpenStudy (lilg132):

well to check which answer is correct you need to remember store A got $250,000 more sales then store B. which would mean 299,862,500 + 250,000 = $300,112,500 (As you can see thats not the same as your answer for store A which means it is wrong) or have i misunderstood the question?

OpenStudy (lilg132):

?

OpenStudy (anonymous):

You have not misunderstood the question, Im on your side with this problem by all means. I was just allowing others to see what I was going through when I solved the problem. I got the same answer as everyone else, the correct answer tht was on the review threw me off.

OpenStudy (lilg132):

ok i get it now lol